# LLM4Mat-Bench: Benchmarking Large Language Models for Materials Property Prediction

### v2

@@ 1 Introduction
With the remarkable success of large language models (LLMs) in solving natural language tasks (Devlin et al., 2018; Radford et al., 2018; Raffel et al., 2020; Achiam et al., 2023; Touvron et al., 2023) and different scientific tasks (Lin et al., 2022; Edwards et al., 2022; Valentini et al., 2023; Castro Nascimento and Pimentel, 2023; Fang et al., 2023; Lv et al., 2024), scientists have recently started to leverage LLMs to tackle very important and challenging problems in materials science, including predicting materials properties (Rubungo et al., 2023; Korolev and Protsenko, 2023; Xie et al., 2023; Das et al., 2023; Qu et al., 2024; Choudhary, 2024) and discovering new materials (Antunes et al., 2023; Gruver et al., 2023; Qu et al., 2024; Choudhary, 2024).
@@
In this work, we introduce LLM4Mat-Bench, a benchmark dataset collected to evaluate the performance of LLMs in predicting the properties of crystalline materials. To the best of our knowledge, LLM4Mat-Bench is the most extensive benchmark to date for assessing the efficacy of language models in materials property prediction. The dataset comprises approximately two million samples, sourced from ten publicly available materials sources, each containing between 10K and 1M structure samples. LLM4Mat-Bench encompasses several tasks, including the prediction of electronic, elastic, and thermodynamic properties based on a material’s composition, crystal information file (CIF), or textual description of its structure. We use LLM4Mat-Bench to evaluate several LLMs of different sizes, namely LLM-Prop (Rubungo et al., 2023) (35M parameters), MatBERT (Walker et al., 2021) (109.5M parameters), and Llama2 (Touvron et al., 2023) (7B parameters). And we provide fixed train-valid-test splits, along with carefully designed zero-shot and few-shot prompts to ensure reproducibility. We anticipate that LLM4Mat-Bench will significantly advance the application of LLMs in addressing critical challenges in materials science, including property prediction and materials discovery.
@@ 2.1.1 Data sources
hMOF (Wilmer et al., 2012) is a publicly available database [^2] https://mof.tech.northwestern.edu/ consisting of about 160K Metal-Organic Frameworks (MOFs), generated by Wilmer et al. using computational approaches. Materials Project (MP) (Jain et al., 2013) is a database with around 150K materials, offering free API access [^3] https://next-gen.materialsproject.org/api for data retrieval, including CIF files and material properties. The Open Quantum Materials Database (OQMD) (Kirklin et al., 2015) is a publicly accessible database [^4] https://www.oqmd.org/ of 1.2M materials, containing DFT-calculated thermodynamic and structural properties, created at Northwestern University. OMDB (Borysov et al., 2017) is an organic materials database with about 12K structures and related electronic band structure properties, freely available [^5] https://omdb.mathub.io/. JARVIS-DFT (Choudhary et al., 2017; Choudhary et al., 2018) is a repository created by NIST researchers, containing around 75.9K material structures with downloadable properties [^6] https://jarvis.nist.gov/jarvisdft. QMOF (Rosen et al., 2021, 2022) is a quantum-chemical property database of over 16K MOFs, accessible via GitHub [^7] https://github.com/Andrew-S-Rosen/QMOF. JARVIS-QETB (Garrity and Choudhary, 2023) is a NIST-created database [^8] https://jarvis.nist.gov/jarvisqetb/ of nearly one million materials with tight-binding parameters for 65 elements. GNoME is a database of 381K new stable materials discovered by Merchant et al. (2023) using graph networks and DFT, available on GitHub [^9] https://github.com/google-deepmind/materials_discovery/blob/main/DATASET.md. Cantor HEA (Li et al., 2024) is a DFT dataset of formation energies for 84K alloy structures, available on Zenodo [^10] https://doi.org/10.5281/zenodo.10854500. SNUMAT is a database with around 10K experimentally synthesized materials and DFT properties, accessible via API [^11] https://www.snumat.com/apis.
@@ 2.1.3 Generating the textual description of crystal structure
Cantor HEA (Li et al., 2024) 84,024 84,019 67,215 8,402 8,402 84K 11M 251M 9.5 868.4 4988.6 4
JARVIS-DFT (Choudhary et al., 2017; Choudhary et al., 2018) 75,965 75,965 60,772 7,596 7,597 76K 9M 25M 5.0 786.0 455.9 20
@@ 3.1 Experimental Details
We conducted about 845 experiments, evaluating the performance of five models and three material representations on each property for each data source. Consistent with standard practices in materials science, we evaluated performance separately for each data source rather than combining samples from different sources for the same property. This approach accounts for variations in techniques and settings used by different data sources, which can result in discrepancies, such as differing band gaps for the same material. Below, we will describe each material representation, model, and metric that we used to conduct our experiments.
@@ 3.1.2 Models
We benchmarked different LLM-based models with various sizes, and a GNN-based baseline. Herein, We provide the details of each model.
CGCNN (Xie and Grossman, 2018) is employed as a GNN baseline which is widely used in the materials science community [^16] Although CGCNN is not state-of-the-art for some properties, it was faster compared to models like ALIGNN (Choudhary and DeCost, 2021) and DeeperGatGNN (Omee et al., 2022), making it suitable for our extensive experiments. We trained on LLM4Mat-Bench from scratch with optimal hyperparameters: 128 hidden dimensions, batch size of 256, three message passing layers, 1e-2 learning rate, 8.0 radius cutoff, 12 nearest neighbors, and 500 training epochs, though extending to 1000 epochs improved performance in some cases.
@@
Llama 2-7b-chat To assess the performance of conversational LLMs in materials property prediction, we tested Llama 2-7b-chat (7 billion parameters) using our designed zero-shot and five-shot prompts (see Table 3) without fine-tuning. For the CIF structure prompts, we removed "# generated using pymatgen" comment that is appended to each file. The maximum input length was set to 4000 tokens while the output length was set to 256, with a batch size of 256 samples, temperature of 0.8, and top-K sampling applied with $K=10$. The details of other models that we compared with Llama 2 can be found in Appendix C. For five-shot examples, we sampled from crystals with shorter structures and descriptions to reduce the context length. We also made sure the property values for those examples are diverse (for instance, they should not all have 0.0 eV as their bandgap values).
@@ 3.1.3 Evaluation Metrics
Table 4: The Wtd. Avg. (MAD/MAE) scores (the higher the better) for the regression tasks in the LLM4Mat-Bench are reported. Bolded results indicate the best model for each input format, while bolded results with blue background highlight the best model per each dataset. Inval. denotes cases where the Llama model failed to generate outputs with a property value or had fewer than 10 valid predictions.
@@
Figure 1: The performance comparison across models for each material representation is presented. The left y-axis shows the log-normalized performance of each LLM-based model relative to the baseline (CGCNN), while the right y-axis (bar plots) displays the average subword tokens per sample for each dataset. Datasets on the x-axis are ordered by increasing average subword tokens. Results for Llama 2-7b-chat:0S and Llama 2-7b-chat:5S are missing in plots (a) and (b), respectively, due to invalid outputs. Higher values in the line plots indicate better performance.
Figure 2: The performance comparison across material representations for each LLM-based model is shown. The y-axis represents the log-normalized Weighted Average (MAD/MAE) score for each representation, while the x-axis displays randomly ordered datasets. In the (a) and (b) plots, some Composition and Structure performance results are missing due to invalid outputs. A higher y-axis value indicates better performance.
@@ 3.2 Discussion
Small, task-specific predictive LLMs exhibit significantly better performance than larger, generative general-purpose LLMs. This performance disparity is evident across both regression (Table 4 and Figure 1) and classification tasks (Table 5) on all 10 datasets. Specifically, LLM-Prop and MatBERT outperform Llama 2-7b-chat:0S and Llama 2-7b-chat:5S by a substantial margin, despite being approximately 200 and 64 times smaller in size, respectively. In regression tasks, LLM-Prop achieves the highest accuracy on 8 out of 10 datasets, with MatBERT leading on the remaining 2 datasets. For classification tasks, both LLM-Prop and MatBERT deliver the best performance on 1 out of 2 datasets. LLM-Prop surpasses MatBERT by 1.8% on the SNUMAT dataset, whereas MatBERT outperforms LLM-Prop by 0.8% on the other dataset. As anticipated, a modest enhancement in average performance is observed across various datasets and input formats when the Llama 2-7b-chat model is evaluated using 5-shot prompts rather than 0-shot prompts. Determining the optimal number of examples required to achieve peak performance will be the focus of future work.
General-purpose generative LLMs hallucinate and often fail to generate valid property values. As shown in Table 4, Table 5, Figure 2, and Appendix E, Llama 2-7b-chat model produces invalid outputs on multiple tasks, where the expected property value is missing. This issue occurs less frequently when the input is a description or chemical formula, but more commonly when the input is a CIF file. One reason may be that descriptions and chemical formulas resemble natural language, which LLMs can more easily interpret compared to CIF files. This may also indicate that when the input modality during inference differs significantly from the modalities encountered during pretraining, fine-tuning is necessary to achieve reasonable performance. Another key observation is that Llama 2-7b-chat model often generates the same property value for different inputs (i.e. hallucinates), contributing to its poor performance across multiple tasks. These findings highlight the importance of caution when using general-purpose generative LLMs for materials property prediction and emphasize the need for fine-tuned, task-specific LLM-based models.
@@
More advanced, general-purpose generative LLMs do not necessarily yield better results in predicting material properties. In Figure 3, we compare the performance of Llama 2-7b-chat-hf model with advanced versions of Llama of comparable sizes when predicting material’s band gap and its stability. Similar comparisons are also conducted for the Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024) models. The results indicate that, despite being trained on substantially larger and higher-quality datasets, more advanced versions of generative LLMs show limited improvements in performance and validity of predictions for material properties. For instance, Llama 3 and 3.1 8b models were trained on over 15 trillion tokens—around eight times more data than the 2 trillion tokens used for the Llama 2 7b models. This finding highlights the ongoing challenges of leveraging LLMs in material property prediction and underscores the need for further research to harness the potential of these robust models in this domain.
@@ E.2 Detailed MAE Results
Input Model FEPA $\downarrow$ Bandgap $\downarrow$ EPA $\downarrow$ Ehull $\downarrow$ Efermi $\downarrow$ Density $\downarrow$ Density atomic $\downarrow$ Volume $\downarrow$ Is stable $\uparrow$ Is gab direct $\uparrow$ Wtd. Avg. (MAD/MAE) $\uparrow$ Wtd. Avg. AUC $\uparrow$
@@
Structure CGCNN (baseline) 0.123 0.366 0.209 0.058 0.607 0.246 1.986 201.186 0.882 0.810 5.319 0.846
Composition Llama 2-7b-chat:0S 128.007 1.913 163.899 195.908 673.939 2.231 15.503 458.456 0.500 0.482 0.389 0.491
@@
Description Llama 2-7b-chat:0S 53.083 1.881 66.971 313.255 288.275 1.647 16.874 424.737 0.500 0.500 0.437 0.500
@@
Input Model FEPA $\downarrow$ Bandgap (OPT) $\downarrow$ Total energy $\downarrow$ Ehull $\downarrow$ Bandgap (MBJ) $\downarrow$ Kv $\downarrow$ Gv $\downarrow$ SLME $\downarrow$ Spillage $\downarrow$ $\epsilon_{x}$ (OPT) $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
@@
Structure CGCNN (baseline) 0.063 0.200 0.078 0.170 0.410 14.470 11.750 5.660 0.400 27.170 7.048
Composition Llama 2-7b-chat:0S 40.722 86.989 91.114 53.109 2.005 125.545 73.335 71.208 Inval. 365.525 Inval.
@@
Description Llama 2-7b-chat:0S 117.691 87.327 90.615 72.709 1.962 108.062 71.895 150.412 3.753 101.906 0.247
@@
Structure CGCNN (baseline) 38.780 34.710 0.190 24.700 50.000 0.120 49.920 442.300 42.420 440.26 7.048
Composition Llama 2-7b-chat:0S 3626.217 120.287 228.849 312.490 109.464 7.708 127.367 868.257 173.183 800.924 Inval.
@@
Description Llama 2-7b-chat:0S 543.798 54.605 273.968 320.801 147.706 11.616 144.830 877.009 203.677 843.939 0.247
@@
Input Model Bandgap GGA $\downarrow$ Bandgap HSE $\downarrow$ Bandgap GGA optical $\downarrow$ Bandgap HSE optical $\downarrow$ Is direct $\uparrow$ Is direct HSE $\uparrow$ SOC $\uparrow$ Wtd. Avg. (MAD/MAE) $\uparrow$ Wtd. Avg. AUC $\uparrow$
@@
Structure CGCNN (baseline) 0.422 0.484 0.668 0.732 0.691 0.675 0.800 1.973 0.722
Composition Llama 2-7b-chat:0S 1.099 1.152 0.998 1.564 0.503 0.484 Inval. 0.940 Inval.
@@
Description Llama 2-7b-chat:0S 1.092 1.161 1.139 1.725 0.499 0.509 Inval. 0.883 Inval.
@@
Input Model FEPA $\downarrow$ Bandgap $\downarrow$ DEPA $\downarrow$ Total energy $\downarrow$ Volume $\downarrow$ Density $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
@@
Structure CGCNN (baseline) 0.014 0.045 0.012 8.431 24.274 0.037 19.478
Composition Llama 2-7b-chat:0S 251.420 2.170 233.95 713.047 425.052 5.633 0.164
@@
Description Llama 2-7b-chat:0S 259.528 3.360 311.944 94.968 295.904 2.576 0.336

### v3

@@ 1 Introduction
With the remarkable success of large language models (LLMs) in solving natural language tasks (Devlin et al., 2018; Radford et al., 2018; Raffel et al., 2020; Achiam et al., 2023; Touvron et al., 2023) and different scientific tasks (Lin et al., 2022; Edwards et al., 2022; Valentini et al., 2023; Castro Nascimento and Pimentel, 2023; Fang et al., 2023; Lv et al., 2024), scientists have recently started to leverage LLMs to tackle very important and challenging problems in materials science, including predicting materials properties (Rubungo et al., 2023; Korolev and Protsenko, 2023; Xie et al., 2023; Das et al., 2023; Qu et al., 2024; Choudhary, 2024; Li et al., 2024b) and discovering new materials (Antunes et al., 2023; Gruver et al., 2023; Qu et al., 2024; Choudhary, 2024).
@@
In this work, we introduce LLM4Mat-Bench, a benchmark dataset collected to evaluate the performance of LLMs in predicting the properties of crystalline materials. To the best of our knowledge, LLM4Mat-Bench is the most extensive benchmark to date for assessing the efficacy of language models in materials property prediction. The dataset comprises approximately two million samples, sourced from ten publicly available materials sources, each containing between 10K and 1M structure samples. LLM4Mat-Bench encompasses several tasks, including the prediction of electronic, elastic, and thermodynamic properties based on a material’s composition, crystal information file (CIF), or textual description of its structure. We use LLM4Mat-Bench to evaluate several LLMs of different sizes, namely LLM-Prop (Rubungo et al., 2023) (35M parameters), MatBERT (Walker et al., 2021) (109.5M parameters), and Llama variants (Touvron et al., 2023; Dubey et al., 2024) (3B, 7B, and 8B parameters), Mistral variants (Jiang et al., 2023) (7B parameters), and Gemma variants (Team et al., 2024a, b) (7B and 9B parameters). And we provide fixed train-valid-test splits, along with carefully designed zero-shot and few-shot prompts to ensure reproducibility. We anticipate that LLM4Mat-Bench will significantly advance the application of LLMs in addressing critical challenges in materials science, including property prediction and materials discovery.
@@ 2.1.1 Data sources
hMOF (Wilmer et al., 2012) is a publicly available database [^2] https://mof.tech.northwestern.edu/ consisting of about 160K Metal-Organic Frameworks (MOFs), generated by Wilmer et al. using computational approaches. Materials Project (MP) (Jain et al., 2013) is a database with around 150K materials, offering free API access [^3] https://next-gen.materialsproject.org/api for data retrieval, including CIF files and material properties. The Open Quantum Materials Database (OQMD) (Kirklin et al., 2015) is a publicly accessible database [^4] https://www.oqmd.org/ of 1.2M materials, containing DFT-calculated thermodynamic and structural properties, created at Northwestern University. OMDB (Borysov et al., 2017) is an organic materials database with about 12K structures and related electronic band structure properties, freely available [^5] https://omdb.mathub.io/. JARVIS-DFT (Choudhary et al., 2017, 2018) is a repository created by NIST researchers, containing around 75.9K material structures with downloadable properties [^6] https://jarvis.nist.gov/jarvisdft. QMOF (Rosen et al., 2021, 2022) is a quantum-chemical property database of over 16K MOFs, accessible via GitHub [^7] https://github.com/Andrew-S-Rosen/QMOF. JARVIS-QETB (Garrity and Choudhary, 2023) is a NIST-created database [^8] https://jarvis.nist.gov/jarvisqetb/ of nearly one million materials with tight-binding parameters for 65 elements. GNoME is a database of 381K new stable materials discovered by Merchant et al. (2023) using graph networks and DFT, available on GitHub [^9] https://github.com/google-deepmind/materials_discovery/blob/main/DATASET.md. Cantor HEA (Li et al., 2024a) is a DFT dataset of formation energies for 84K alloy structures, available on Zenodo [^10] https://doi.org/10.5281/zenodo.10854500. SNUMAT is a database with around 10K experimentally synthesized materials and DFT properties, accessible via API [^11] https://www.snumat.com/apis.
@@ 2.1.3 Generating the textual description of crystal structure
Cantor HEA (Li et al., 2024a) 84,024 84,019 67,215 8,402 8,402 84K 11M 251M 9.5 868.4 4988.6 4
JARVIS-DFT (Choudhary et al., 2017, 2018) 75,965 75,965 60,772 7,596 7,597 76K 9M 25M 5.0 786.0 455.9 20
@@ 3.1 Experimental Details
We conducted about 1,235 experiments, evaluating the performance of five models and three material representations on each property for each data source. Consistent with standard practices in materials science, we evaluated performance separately for each data source rather than combining samples from different sources for the same property. This approach accounts for variations in techniques and settings used by different data sources, which can result in discrepancies, such as differing band gaps for the same material. Below, we will describe each material representation, model, and metric that we used to conduct our experiments.
@@ 3.1.2 Models
We benchmarked different LLM-based models with various sizes and types, and a GNN-based baseline. Herein, We provide the details of each model.
CGCNN (Xie and Grossman, 2018) is employed as a GNN baseline which is widely used in the materials science community [^16] Although CGCNN is not state-of-the-art for some properties, it was faster compared to models like ALIGNN (Choudhary and DeCost, 2021) and DeeperGatGNN (Omee et al., 2022), making it suitable for our extensive experiments. We trained it on LLM4Mat-Bench from scratch with optimal hyperparameters: 128 hidden dimensions, batch size of 256, three message passing layers, 1e-2 learning rate, 8.0 radius cutoff, 12 nearest neighbors, and 500 training epochs, though extending to 1000 epochs improved performance in some cases.
@@
Llama, Gemma, and Mistral To assess the performance of conversational LLMs in materials property prediction, we tested the currently available variants of Llama, Gemma, and Mistral using our designed zero-shot and five-shot prompts (see Table 3) without fine-tuning. For the CIF structure prompts, we removed "# generated using pymatgen" comment that is appended to each file. The maximum input length was set to the maximum number of tokens each model can handle while the output length was set to 256, with a batch size of 256 samples, temperature of 0.8, and top-K sampling applied with $K=10$. The inference details of each model can be found in Appendix C. For five-shot examples, we sampled from crystals with shorter structures and descriptions to reduce the context length. We also made sure the property values for those examples are diverse (for instance, they should not all have 0.0 eV as their bandgap values). For each model family, we first compared all the associated models for band gap and stability prediction on the MP test set and found no significant performance gain from one variant to another (see Figure 3). Therefore, for the remaining properties in each dataset, we conducted experiments using only one variant from each family as its representative. Additionally, we reported both zero-shot and five-shot performance for Llama, while focusing solely on five-shot performance for other models due to the significant performance gap observed between zero-shot and five-shot scenarios.
@@ 3.1.3 Evaluation Metrics
Table 4: The Wtd. Avg. (MAD:MAE) scores (the higher the better) for the regression tasks in the LLM4Mat-Bench are reported. Bolded results indicate the best model for each input format, while bolded results with blue background highlight the best model per each dataset. Inval. denotes cases where the Llama model failed to generate outputs with a property value or had fewer than 10 valid predictions.
@@
Mistral 7b-Instruct-v0.1:5S Inval. 0.328 0.174 Inval. 0.095 0.251 0.236 Inval. Inval. Inval.
Gemma 2-9b-it:5S Inval. Inval. 0.642 Inval. 0.8 0.763 0.957 Inval. Inval. Inval.
@@
Mistral 7b-Instruct-v0.1:5S Inval. 0.328 0.134 Inval. 0.147 0.459 0.595 Inval. 0.891 Inval.
Gemma 2-9b-it:5S Inval. Inval. Inval. Inval. 1.013 0.674 0.919 Inval. 1.194 0.626
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval.
Gemma 2-9b-it:5S Inval. 0.522
@@
Figure 1: The performance comparison across models for each material representation is presented. The left y-axis shows the log-normalized performance of each LLM-based model relative to the baseline (CGCNN), while the right y-axis (bar plots) displays the average subword tokens per sample for each dataset. Datasets on the x-axis are ordered by increasing average subword tokens. Results for some chat-like models are missing in each subplot due to invalid outputs on at least one of the property. Higher values in the line plots indicate better performance. Panels (a), (b), and (c) represents the performance comparison where the input is a chemical composition, CIF, and structure description, respectively.
Figure 2: The performance comparison across material representations for each LLM-based model is shown. The y-axis represents the log-normalized Weighted Average (MAD:MAE) score for each representation, while the x-axis displays randomly ordered datasets. In the (a)-(d) plots, some Composition and Structure performance results are missing due to invalid outputs. A higher y-axis value indicates better performance. Panels (a) to (f) represents the results for Llama 2-7b-chat:0S, Llama 2-7b-chat:5S, Mistral 7b-Instruct-v0.1:5S, Gemma 2-9b-it:5S, MatBERT, and LLM-Prop, respectively.
@@ 3.2 Discussion
Small, task-specific predictive LLMs exhibit significantly better performance than larger, generative general-purpose LLMs. This performance disparity is evident across both regression (Table 4 and Figure 1) and classification tasks (Table 5) on all 10 datasets. Specifically, LLM-Prop and MatBERT outperform conversational LLMs by a substantial margin, despite being approximately 200 and 64 times smaller in size, respectively. In regression tasks, LLM-Prop achieves the highest accuracy on 8 out of 10 datasets, with MatBERT leading on the remaining 2 datasets. For classification tasks, both LLM-Prop and MatBERT deliver the best performance on 1 out of 2 datasets. LLM-Prop surpasses MatBERT by 1.8% on the SNUMAT dataset, whereas MatBERT outperforms LLM-Prop by 0.8% on the other dataset. As anticipated, a modest enhancement in average performance is observed across various datasets and input formats when the Llama 2-7b-chat model is evaluated using 5-shot prompts rather than 0-shot prompts. Determining the optimal number of examples required to achieve peak performance will be the focus of future work.
General-purpose generative LLMs hallucinate and often fail to generate valid property values. As shown in Table 4, Table 5, Figure 2, and Appendix E, Llama, Gemma, and Mistral models produce invalid outputs on multiple tasks, where the expected property value is missing. This issue occurs less frequently when the input is a description or chemical formula, but more commonly when the input is a CIF file. One reason may be that descriptions and chemical formulas resemble natural language, which LLMs can more easily interpret compared to CIF files. This may also indicate that when the input modality during inference differs significantly from the modalities encountered during pretraining, fine-tuning is necessary to achieve reasonable performance. Another key observation is that these models often generate the same property value for different inputs (i.e. hallucinate), contributing to their poor performance across multiple tasks. These findings highlight the importance of caution when using general-purpose generative LLMs for materials property prediction and emphasize the need for fine-tuned, task-specific LLM-based models.
@@
More advanced, general-purpose generative LLMs do not necessarily yield better results in predicting material properties. In Figure 3, we compare the performance of Llama 2-7b-chat-hf model with advanced versions of Llama of comparable sizes when predicting material’s band gap and its stability. Similar comparisons are also conducted for the Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024a) models. The results indicate that, despite being trained on substantially larger and higher-quality datasets, more advanced versions of generative LLMs show limited improvements in performance and validity of predictions for material properties. For instance, Llama 3 and 3.1 8b models were trained on over 15 trillion tokens—around eight times more data than the 2 trillion tokens used for the Llama 2 7b models. This finding highlights the ongoing challenges of leveraging LLMs in material property prediction and underscores the need for further research to harness the potential of these robust models in this domain.
@@ E.1 Detailed MAD:MAE Results
Mistral 7b-Instruct-v0.1:5S 0.025 Inval. 0.063 0.071 0.249 0.107 0.369 Inval. Inval. Inval.
Gemma 2-9b-it:5S 0.696 1.181 0.793 0.777 0.834 0.762 0.953 Inval. 0.497 Inval.
@@
Mistral 7b-Instruct-v0.1:5S 0.126 0.951 0.017 0.154 Inval. Inval. Inval. 0.866 0.499 0.500
Gemma 2-9b-it:5S Inval. 0.946 Inval. Inval. 0.814 Inval. Inval. Inval. Inval. Inval.
@@
Mistral 7b-Instruct-v0.1:5S 1.484 1.315 0.011 0.575 0.947 0.216 0.49 0.221 0.813 0.661
Gemma 2-9b-it:5S 1.768 0.445 0.695 Inval. Inval. 0.366 0.609 Inval. 0.982 Inval.
@@
Mistral 7b-Instruct-v0.1:5S 1.484 1.315 0.011 0.575 0.947 0.216 0.49 0.221 0.813 0.661
Gemma 2-9b-it:5S 1.576 1.901 1.131 Inval. Inval. 0.439 0.57 Inval. Inval. Inval.
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval. Inval. 0.738 0.494 0.500 Inval.
Gemma 2-9b-it:5S Inval. Inval. 0.991 0.941 0.5 0.607 0.5
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval. Inval. Inval. 0.005 0.565 Inval.
Gemma 2-9b-it:5S 0.688 Inval. Inval. Inval. Inval. Inval. Inval.
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval. Inval. Inval. 0.015 0.563 0.366
Gemma 2-9b-it:5S 0.701 Inval. Inval. Inval. Inval. Inval. Inval.
@@
Mistral 7b-Instruct-v0.1:5S 0.125 0.068 0.377 0.018
Gemma 2-9b-it:5S 1.013 Inval. Inval. Inval.
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval. Inval. Inval.
Gemma 2-9b-it:5S Inval. 0.566 Inval. Inval.
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval. Inval. Inval.
Gemma 2-9b-it:5S Inval. 0.615 Inval. Inval.
@@ E.2 Detailed MAE Results
Input Model FEPA $\downarrow$ Bandgap $\downarrow$ EPA $\downarrow$ Ehull $\downarrow$ Efermi $\downarrow$ Density $\downarrow$ Density atomic $\downarrow$ Volume $\downarrow$ Is stable $\uparrow$ Is gab direct $\uparrow$ Wtd. Avg. Wtd. Avg.
(MAD:MAE) $\uparrow$ ROC AUC $\uparrow$
@@
CIF CGCNN (baseline) 0.123 0.366 0.209 0.058 0.607 0.246 1.986 201.186 0.882 0.810 5.319 0.846
Comp. Llama 2-7b-chat:0S 128.007 1.913 163.899 195.908 673.939 2.231 15.503 458.456 0.500 0.482 0.389 0.491
@@
Mistral 7b-Instruct-v0.1:5S 40.834 Inval. 23.89 3.182 8.992 20.078 31.709 Inval. Inval. Inval. Inval. Inval.
Gemma 2-9b-it:5S 1.44 1.009 1.905 0.289 2.685 2.832 12.27 Inval. 0.497 Inval. Inval. Inval.
@@
Descr. Llama 2-7b-chat:0S 53.083 1.881 66.971 313.255 288.275 1.647 16.874 424.737 0.500 0.500 0.437 0.500
@@
Mistral 7b-Instruct-v0.1:5S 7.955 1.252 88.691 1.461 Inval. Inval. Inval. 395.496 0.499 0.500 Inval. 0.499
Gemma 2-9b-it:5S Inval. 1.259 Inval. Inval. 2.75 Inval. Inval. Inval. Inval. Inval. Inval. Inval.
@@
Input Model FEPA $\downarrow$ Bandgap (OPT) $\downarrow$ Total energy $\downarrow$ Ehull $\downarrow$ Bandgap (MBJ) $\downarrow$ Kv $\downarrow$ Gv $\downarrow$ SLME $\downarrow$ Spillage $\downarrow$ $\epsilon_{x}$ (OPT) $\downarrow$ Wtd. Avg. $\uparrow$
(MAD:MAE) $\uparrow$
@@
CIF CGCNN (baseline) 0.063 0.200 0.078 0.170 0.410 14.470 11.750 5.660 0.400 27.170 7.048
Comp. Llama 2-7b-chat:0S 40.722 86.989 91.114 53.109 2.005 125.545 73.335 71.208 Inval. 365.525 Inval.
@@
Descr. Llama 2-7b-chat:0S 117.691 87.327 90.615 72.709 1.962 108.062 71.895 150.412 3.753 101.906 0.247
@@
Mistral 7b-Instruct-v0.1:5S 19.0 86.39 90.845 1.704 2.181 62.534 28.427 18.595 2.067 58.236 0.328
Gemma 2-9b-it:5S 27.569 7.642 0.217 Inval. Inval. 0.491 195.223 Inval. Inval. Inval. Inval.
@@
CIF CGCNN (baseline) 38.780 34.710 0.190 24.700 50.000 0.120 49.920 442.300 42.420 440.26 7.048
Comp. Llama 2-7b-chat:0S 3626.217 120.287 228.849 312.490 109.464 7.708 127.367 868.257 173.183 800.924 Inval.
@@
Mistral 7b-Instruct-v0.1:5S 27.317 11.241 23.747 246.496 56.247 1.182 227.868 2785.193 197.133 1098.064 0.328
Gemma 2-9b-it:5S 24.575 32.609 0.353 Inval. Inval. 0.589 182.664 Inval. 171.171 Inval. Inval.
@@
Descr. Llama 2-7b-chat:0S 543.798 54.605 273.968 320.801 147.706 11.616 144.830 877.009 203.677 843.939 0.247
@@
Mistral 7b-Instruct-v0.1:5S 29.283 11.046 23.322 76.812 44.446 0.999 227.239 3155.212 206.891 1058.133 0.328
Gemma 2-9b-it:5S 1.483 87.004 90.374 0.937 2.327 Inval. Inval. 21.83 0.805 55.716 Inval.
@@
Input Model Bandgap GGA $\downarrow$ Bandgap HSE $\downarrow$ Bandgap GGA optical $\downarrow$ Bandgap HSE optical $\downarrow$ Is direct $\uparrow$ Is direct HSE $\uparrow$ SOC $\uparrow$ Wtd. Avg Wtd. Avg.
(MAD:MAE) $\uparrow$ ROC AUC $\uparrow$
@@
CIF CGCNN (baseline) 0.422 0.484 0.668 0.732 0.691 0.675 0.800 1.973 0.722
Comp. Llama 2-7b-chat:0S 1.099 1.152 0.998 1.564 0.503 0.484 Inval. 0.940 Inval.
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval. Inval. 1.821 0.404 0.500 Inval. Inval. Inval.
Gemma 2-9b-it:5S Inval. Inval. 1.164 1.427 0.500 0.607 0.500 Inval. 0.522
@@
Descr. Llama 2-7b-chat:0S 1.092 1.161 1.139 1.725 0.499 0.509 Inval. 0.883 Inval.
@@
Input Model FEPA $\downarrow$ Bandgap $\downarrow$ DEPA $\downarrow$ Total energy $\downarrow$ Volume $\downarrow$ Density $\downarrow$ Wtd. Avg.
(MAD:MAE) $\uparrow$
@@
CIF CGCNN (baseline) 0.014 0.045 0.012 8.431 24.274 0.037 19.478
Comp. Llama 2-7b-chat:0S 251.420 2.170 233.95 713.047 425.052 5.633 0.164
@@
Descr. Llama 2-7b-chat:0S 259.528 3.360 311.944 94.968 295.904 2.576 0.336